Assuming the ideal C/a ratio for the HCP Ti (with a=2.95 Angstrum ), the radius of the largest sphere that will fit interstitially in Ti is?
1
Expert's answer
2020-06-16T10:41:54-0400

As per the question,

Question is incomplete, here c/a ratio is not given let it is k

"\frac{c}{a}=k"

"a=2.95A^\circ"

"h=\sqrt{a^2-\frac{a}{2}^2}=\sqrt{\frac{3}{4}}a"

x"=\frac{2}{3}h=\frac{2}{3}\sqrt{\frac{3}{4}}a =\frac{a}{\sqrt{3}}"